OpenStudy (daisy.xoxo619):

The constant rate of growth for the function f(x)=76,042*1.148^x is _________%

OpenStudy (campbell_st):

ok... so the compound interest formula... which is basically what your problem is similar to is \[A = P \times (1 + \frac{r}{100})^n\] in your equation P = 76042 in my equation r is the growth rate so comparing you have \[1.148 = 1 + \frac{r}{100}\] so what do you think r is..?

OpenStudy (daisy.xoxo619):

is r=114.8?

OpenStudy (daisy.xoxo619):

@ganeshie8 @mathmale @mathstudent55

OpenStudy (campbell_st):

well not quite 1 + r = 1.148 so subtract 1 from both sides r = 0.148 now convert it to a percentage